### v1

# Carrier Aggregation Enabled MIMO-OFDM Integrated Sensing and Communication

Abstract
In the evolution towards the forthcoming era of sixth-generation (6G) mobile communication systems characterized by ubiquitous intelligence, integrated sensing and communication (ISAC) is in a phase of burgeoning development. However, the capabilities of communication and sensing within single frequency band fall short of meeting the escalating demands. To this end, this paper introduces a carrier aggregation (CA)-enabled multi-input multi-output orthogonal frequency division multiplexing (MIMO-OFDM) ISAC system fusing the sensing data on high and low-frequency bands by symbol-level fusion for ultimate communication experience and high-accuracy sensing. The challenges in sensing signal processing introduced by CA include the initial phase misalignment of the echo signals on high and low-frequency bands due to attenuation and radar cross section, and the fusion of the sensing data on high and low-frequency bands with different physical-layer parameters. To this end, the sensing signal processing is decomposed into two stages. In the first stage, the problem of initial phase misalignment of the echo signals on high and low-frequency bands is solved by the angle compensation, space-domain diversity and vector cross-correlation operations. In the second stage, this paper realizes symbol-level fusion of the sensing data on high and low-frequency bands through sensing vector rearrangement and cyclic prefix adjustment operations, thereby obtaining high-precision sensing performance. Then, the closed-form communication mutual information (MI) and sensing Cramér-Rao lower bound (CRLB) for the proposed ISAC system are derived to explore the theoretical performance bound with CA. Simulation results validate the feasibility and superiority of the proposed ISAC system.
Index Terms:
Carrier aggregation (CA), integrated sensing and communication (ISAC), multi-band cooperative ISAC, multi-input multi-output (MIMO), orthogonal frequency division multiplexing (OFDM).
I Introduction
The next-generation mobile communication system is anticipated to offer a plethora of multi-dimensional services extending beyond mere communication, such as smart transportation, smart factories, and digital twins [1], which have raised the demands for high-speed communication and high-accuracy sensing. To this end, integrated sensing and communication (ISAC) is considered to provide the high-speed communication experience and high-accuracy sensing [2, 3, 4, 5].
Nevertheless, the limitations imposed by sensing under single frequency band have spurred the study of ISAC over multiple frequency bands with the technique of carrier aggregation (CA). CA-enabled ISAC systems have the following potential advantages.
  • Communication: CA enhances bandwidth utilization and boosts communication transmission data rate while supporting various wireless access technologies, thereby improving overall communication performance [6].
  • Sensing: CA empowers sensing systems with expanded bandwidth, which facilitates high-resolution sensing and anti-noise performance [7, 5].
Therefore, it is reasonable to utilize CA to empower ISAC mobile communication system, which is expected to bring expeditious communication and precision sensing.
TABLE I: A summary for the related work, with the abbreviations OFDM: orthogonal frequency division multiplexing, DFT: discrete Fourier transmit, CS: compressed sensing, NC: non-continuous, BWE: bandwidth extrapolation.
Types
related
work
Innovation Shortcoming
Intra-band CA [8]
Introducing the concept of steeped-carrier OFDM
radar to achieve high-resolution sensing
Only suitable for low speed targets
[9]
Proposing a novel DFT method to address the
shortcoming in [8]
Poor anti-noise performance
[10]
An ISAC system under NC fragmented spectrum bands
is proposed to enhance spectrum utilization
The degradation of Fourier sidelobes
[11]
A CS-based sensing method is proposed
to overcome the shortcoming in [10]
High computational complexity
Inter-band CA [12]
A BWE method is proposed to recovery
the data on missing spectrum bands
The resolution performance is limited
[13]
Extending the BWE method to polarimetric radar
results in a higher enhancement in resolution
Unsuitable for OFDM signal
[7]
Joint high and low-frequency bands OFDM signal
for high-accuracy sensing
Unsuitable for multiple antennas system
The researches on CA-enabled ISAC systems are classified into two types: intra-band CA and inter-band CA. The main challenges of CA-enabled ISAC system lie in signal design and processing. A summary of the related work is shown in Table I and the details are as follows.
  • Intra-band CA: Two primary ISAC signals are designed for intra-band CA: steeped-carrier OFDM ISAC signal and non-continuous (NC) fragmented OFDM ISAC signals. In terms of steeped-carrier OFDM ISAC signals, Pfeffer et al. [8] proposed the concept of steeped-carrier OFDM signals, which assigns different carrier frequencies to each resource block group during the continuous OFDM symbol times, expanding the overall bandwidth and enhancing the sensing performance. However, it fails to consider the Doppler frequency shift induced by rapidly moving targets. To this end, Schweizer et al. [9] explored a modified DFT method to mitigate the phase error caused by the high mobility of targets in steeped-carrier OFDM scheme, which obtains a high-accuracy velocity estimation. In terms of NC fragmented OFDM ISAC signals, Huang et al. [10] introduced a NC fragmented OFDM ISAC system, which can flexibly allocate spectrum bands and adapt to dynamic spectrum environment. However, the missing spectrum bands deteriorate Fourier sidelobes in target sensing. To address this problem, Liu et al. [11] proposed a joint compressed sensing (CS) and machine learning sensing method for fragmented spectrum bands to achieve high-accuracy estimation of range and velocity with low sidelobes.
  • Inter-band CA: Multi-band radar stands as a representative example in inter-band CA. Cuomo [12] discussed a NC multi-band coherent radar and proposed a bandwidth extrapolation (BWE) method to recover missing spectrum bands, thereby enhancing the range resolution of radar. Suwa et al. [13] extended the BWE method to polarimetric radar, which achieves high resolution of range estimation.
In summary, intra-band and inter-band CA enhance the efficient utilization of fragmented spectrum bands, leading to performance gains in both communication and sensing. Currently, the fragmented spectrum bands in mobile communication system are mostly dispersed, indicating broader applications in mobile communication system under inter-band CA. However, the investigation into inter-band CA in ISAC mobile communication systems is notably scarce. In our previous work, we have investigated the CA-enabled OFDM ISAC system, corroborating the effectiveness of CA through the validation [7]. However, our initial exploration, although valuable, remains somewhat rudimentary and idealized, overlooking the typical MIMO-OFDM signal in mobile communication systems.
Hence, this paper investigates a CA-enabled MIMO-OFDM ISAC system where the fragmented high and low-frequency bands are shared for downlink (DL) communication and potential target sensing. In addition, a CA-enabled MIMO-OFDM ISAC signal model and the corresponding sensing processing method are presented. Through theoretical derivations and numerical simulations, we validate the superiority of the proposed ISAC system over conventional ISAC systems. The main contributions of this paper are summarized as follows.
  • We present a CA-enabled MIMO-OFDM ISAC system aiming at achieving unparalleled communication experience and high-accuracy sensing. By introducing CA and MIMO, the communication capacity of ISAC system is significantly enhanced, as well as improving the accuracy and anti-noise capability of sensing. The simulation results validate the feasibility and efficiency of the proposed ISAC system, demonstrating its advantages over conventional ISAC systems.
  • In terms of sensing processing, we have delved into the data-level and symbol-level fusion of the sensing data on high and low-frequency bands. In the symbol-level fusion of sensing data, two challenges manifest: 1) The initial phase misalignment of the echo signals on high and low-frequency bands due to attenuation and radar cross section (RCS); 2) The fusion of the sensing data on high and low-frequency bands with different physical-layer parameters. To this end, this paper decomposes the sensing signal processing into signal preprocessing stage and sensing information fusion stage. The signal preprocessing stage fully leverage the sensing data from multiple antennas while aligning the initial phases of the echo signals on high and low-frequency bands using the cyclic cross-correlation (CCC) method, as detailed in Section III-A. In the sensing information fusion stage, this paper realizes the symbol-level fusion of the sensing data on high and low-frequency bands, as well as the range and velocity estimations of target by rearranging the high and low-frequency feature vectors and using the premise of cyclic prefix (CP) dynamic adjustment, as detailed in Section III-B.
  • The theoretical boundaries of the proposed CA-enabled MIMO-OFDM ISAC signal are derived. Specifically, mutual information (MI) is adopted as the performance metric for communication, while the Cramér-Rao lower bound (CRLB) is utilized as a typical performance metric of sensing. Based on the proposed ISAC system, we derived the closed-form communication MI under the frequency-selective fading MIMO-OFDM channel model and sensing CRLBs for range and velocity estimations under the MIMO-OFDM sensing signal model.
The rest of the paper is organized as follows. Section II introduces the system model, including the communication model and sensing model. In Section III, a novel CA-enabled MIMO-OFDM ISAC signal processing method is proposed. Section IV derives the communication MI and sensing CRLB of the proposed CA-enabled MIMO-OFDM ISAC signal. The simulation results are provided in Section V, while the conclusions are outlined in Section VI.
[…]
II System Model
We consider a CA-enabled MIMO-OFDM ISAC base station (BS) for DL communication and target sensing as shown in Fig. 1. The ISAC BS is equipped with $N_{\text{T}}+N_{\text{R}}$ antennas, where $N_{\text{T}}$ transmit antennas in transmitter (Tx) simultaneously serve $U$ user equipments (UEs) with $N_{\text{U}}$ antennas and detect potential targets. $N_{\text{R}}$ receive antennas in receiver (Rx) are equipped to receive the echo signals reflected by targets [14]. The spacing of antennas in BS is denoted by $d_{\text{r}}$ [15].
Figure 1: CA-enabled MIMO-OFDM ISAC system.
II-A CA-enabled MIMO-OFDM ISAC Signal
For the $b\in{\mathcal{B}}=\{1,2\}$-th carrier component (CC) of the ISAC BS, a total number of $N_{b}$ subcarriers and $M_{b}$ OFDM symbols are used for DL communication and sensing, where $n\in\{0,1,\cdots,N_{b}-1\}$ and $m\in\{0,1,\cdots,M_{b}-1\}$ denote the indices of subcarriers and OFDM symbols, respectively. For the $b$-th CC, the symbol data matrix [16]
$\mathbf{S}^{b}=\left[\mathbf{S}_{0}^{b},\mathbf{S}_{1}^{b},\cdots,\mathbf{S}_{N_{b}-1}^{b}\right]^{\text{T}}\in\mathbb{C}^{UN_{b}\times M_{b}}$ (1)
represents the data for all UEs on $N_{b}$ subcarriers during $M_{b}$ OFDM symbol times, where $\mathbf{S}_{n}^{b}=\left[\mathbf{s}_{n,1}^{b},\mathbf{s}_{n,2}^{b},\cdots,\mathbf{s}_{n,U}^{b}\right]\in\mathbb{C}^{M_{b}\times U}$ indicates the data for all UEs on the $n$-th subcarrier during $M_{b}$ symbol times with $\mathbf{s}_{n,u}^{b}\in\mathbb{C}^{M_{b}\times 1}$ being the specific symbol data vector for the $u\in\{1,2,\cdots,U\}$-th UE [16]. To guarantee that the communication beam focuses on UEs and potential targets, the symbol data matrix undergoes a precoding operation [17]. Specifically, a compact precoding matrix [16]
$\mathbf{W}^{b}=\mathrm{diag}\left(\mathbf{W}_{0}^{b},\mathbf{W}_{1}^{b},\cdots,\mathbf{W}_{N_{b}-1}^{b}\right)\in\mathbb{C}^{N_{b}N_{\text{T}}\times N_{b}U}$ (2)
is designed for all UEs on $N_{b}$ subcarriers. $\mathbf{W}_{n}^{b}=\left[\mathbf{w}_{n,1}^{b},\mathbf{w}_{n,2}^{b},\cdots,\mathbf{w}_{n,U}^{b}\right]\in\mathbb{C}^{N_{\text{T}}\times U}$ denotes the precoding matrix for all UEs on the $n$-th subcarriers with $\mathbf{w}_{n,u}^{b}\in\mathbb{C}^{N_{\text{T}}\times 1}$ being the specific precoding vector for the $u$-th UE [16]. Therefore, the transmit precoded data on $N_{b}$ subcarriers during $M_{b}$ symbol times is [16]
$\mathbf{X}^{b}$ $=\mathbf{W}^{b}\mathbf{S}^{b}$ (3)
$=\left[(\mathbf{X}_{0}^{b})^{\text{T}},(\mathbf{X}_{1}^{b})^{\text{T}},\cdots,(\mathbf{X}_{N_{b}-1}^{b})^{\text{T}}\right]^{\text{T}}\in\mathbb{C}^{N_{b}N_{\text{T}}\times M_{b}},$
where $\mathbf{X}_{n}^{b}=\mathbf{W}^{b}(\mathbf{S}_{n}^{b})^{\text{T}}\in\mathbb{C}^{N_{\text{T}}\times M_{b}}$ is the data on the $n$-th subcarrier during $M_{b}$ OFDM symbol times. The transmit precoded data is transformed into time-domain signal by undergoing an inverse discrete Fourier transform (IDFT).
Therefore, the baseband time-domain transmit signal by the $k$-th transmit antenna on the $n^{\prime}$-th subcarrier during the $m$-th OFDM symbol time in the $b$-th CC is [18]
$x^{b}\left(k,n^{\prime},m\right)=\frac{1}{N_{b}}\sum_{n=0}^{N_{b}-1}\mathbf{X}_{n}^{b}\left(k,m\right)e^{j2\pi\frac{n^{\prime}n}{N_{b}}},$ (4)
where $\mathbf{X}_{n}^{b}\left(k,m\right)$ denotes the $\left(k,m\right)$-th element of $\mathbf{X}_{n}^{b}$, $n^{\prime}\in\{0,1,\cdots,N_{b}-1\}$ represents the index of subcarriers in time-domain and $k\in\{0,1,\cdots,N_{\text{T}}-1\}$ represents the index of transmit antenna. A CP with size $N_{\text{s}}^{b}$ greater than the maximum delay caused by target reflection is added to avoid inter-symbol interference. After digital-to-analog conversion (DAC), the analog signals of the $1$-st and $2$-nd CCs are up-converted using different local oscillators (LOs), mixed, and transmitted via a high power amplifier (HPA).
Overall, the CA-enabled MIMO-OFDM ISAC signal by the $k$-th transmit antenna on the $n$-th subcarrier during the $m$-th OFDM symbol time is expressed as
$x(t)_{k}=\sum_{b=1}^{\mathcal{B}}\sum_{m=0}^{M_{b}-1}\sum_{n=0}^{N_{b}-1}\mathbf{X}_{n}^{b}\left(k,m\right)e^{j2\pi t(f_{\mathrm{C}}^{b}+n\Delta f^{b})}\text{rect}\left(\frac{t-mT^{b}}{T^{b}}\right),$ (5)
where $f_{\mathrm{C}}^{b}$ denotes the carrier frequency of the $b$-th CC; $\Delta f^{b}=1/T_{\text{ofdm}}^{b}$ is the subcarrier spacing of the $b$-th CC, where $T_{\text{ofdm}}^{b}$ is the elementary symbol duration [18]; $T^{b}=T_{\text{ofdm}}^{b}+T_{\text{s}}^{b}$ is the total symbol duration with $T_{\text{s}}^{b}=(N_{\text{s}}^{b}T_{\text{ofdm}}^{b})/N_{b}$ being an adjustable CP duration.
II-B Communication Model
In the communication channel model, the differences in attenuations caused by high and low-frequency bands are not considered.
II-B1 Communication channel model
We utilize the widely adopted multi-path channel model in MIMO channels, assuming the presence of $L$ delay paths in the communication environment. The index of receive antennas in UE is denoted by $z\in\{0,1,\cdots,N_{\text{U}}-1\}$ and
${\mathbf{H}_{u}^{l,b}=\left[\begin{array}[]{cccc}h_{0,0}^{u,l,b}&h_{0,1}^{u,l,b}&\cdots&h_{0,N_{\text{U}}-1}^{u,l,b}\\ h_{1,0}^{u,l,b}&h_{1,1}^{u,l,b}&\cdots&h_{1,N_{\text{U}}-1}^{u,l,b}\\ \vdots&\vdots&\ddots&\vdots\\ h_{N_{\text{T}}-1,0}^{u,l,b}&h_{N_{\text{T}}-1,1}^{u,l,b}&\cdots&h_{N_{\text{T}}-1,N_{\text{U}}-1}^{u,l,b}\end{array}\right]}\in\mathbb{C}^{N_{\text{T}}\times N_{\text{U}}}$ (6)
is the channel matrix of the $l\in\{0,1,\cdots,L-1\}$-th delay path for the $u$-th UE with $h_{k,z}^{u,l,b}\sim\mathcal{CN}(0,1)$ being the corresponding channel element with independent identically distributed (i.i.d.) Rayleigh fading coefficients [16]. Therefore, the MIMO communication channel matrix in time-domain for the $u$-th UE on the $n$-th subcarrier during the $m$-th OFDM symbol time is expressed as [18]
$\mathbf{H}_{u,m}^{n}=\sum_{l=0}^{L-1}\mathbf{H}_{u}^{l,b}\delta{\left[m-l\right]}\in\mathbb{C}^{N_{\text{T}}\times N_{\text{U}}}.$ (7)
II-B2 Received communication signal
At the DL preamble information period, pilot preambles are used to obtain channel state information (CSI) estimation [19]. At the DL data information period, the $u$-th UE receives the DL communication signal, and undergoes down-conversion, analog-to-digital conversion (ADC), CP removal, discrete Fourier transform (DFT), and other operations to get a received data $\mathbf{Y}_{u}^{\mathrm{C}}$, which is expressed as
$\mathbf{Y}_{u}^{\mathrm{C}}=\sum_{b=1}^{\mathcal{B}}\mathbf{X}_{\mathrm{C}}^{b}\mathbf{H}_{\mathrm{C}}^{b}+\mathbf{W}_{u}^{\mathrm{C}},$ (8)
where $\mathbf{H}_{\mathrm{C}}^{b}=\left[(\widetilde{\mathbf{H}}_{u,m}^{0})^{\text{T}},\cdots,(\widetilde{\mathbf{H}}_{u,m}^{N_{b}-1})^{\text{T}}\right]\in\mathbb{C}^{N_{b}N_{\text{T}}\times N_{\text{U}}}$ with $\widetilde{\mathbf{H}}_{u,m}^{n}=\sum_{l=0}^{L-1}\widetilde{\mathbf{H}}_{u}^{l,b}e^{\frac{-j2\pi nl}{N_{b}}}\in\mathbb{C}^{N_{\text{T}}\times N_{\text{U}}}$, while $\widetilde{\mathbf{H}}_{u}^{l,b}$ is the estimated CSI matrix; $\mathbf{X}_{\mathrm{C}}^{b}=\text{diag}\left((\mathbf{X}_{0}^{b})^{\text{T}},\cdots,(\mathbf{X}_{N_{b}-1}^{b})^{\text{T}}\right)\in\mathbb{C}^{N_{b}M_{b}\times N_{b}N_{\text{T}}}$ is the transmit data; $\mathbf{W}_{u}^{\mathrm{C}}\sim\mathcal{CN}\left(0,\sigma_{\mathrm{C}}^{2}\right)$ is additive Gaussian white noise (AWGN) of dual-CC communication channel.
In terms of communication performance in CA-enabled MIMO-OFDM ISAC system, the closed-form communication MI is derived in Section IV-A.
II-C Sensing Model
In the sensing channel model, the differences in attenuations caused by high and low-frequency bands are considered. Then, the sensing channel model of the $b$-th CC and received sensing signal are derived.
II-C1 Sensing channel model
We assume that the radial velocity of a potential target is $v_{0}$ and the range to BS is $r_{0}$. Meanwhile, the angle of arrive (AoA) and the angle of departure (AoD) between the target and BS are identical and are denoted by $\theta_{\text{Rx}}=\theta_{\text{Tx}}$. The transmit and received steering vectors are expressed as (9) and (10), respectively.
$\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})$ $=\left[e^{j2\pi p(\frac{d_{\text{r}}}{\lambda^{b}})\sin(\theta_{\text{Rx}})}\right]^{\text{T}}|_{p=0,1,\cdots,N_{\text{R}}-1},$ (9)
$\mathbf{a}_{\text{Tx}}(\theta_{\text{Tx}})$ $=\left[e^{j2\pi k(\frac{d_{\text{r}}}{\lambda^{b}})\sin(\theta_{\text{Tx}})}\right]^{\text{T}}|_{k=0,1,\cdots,N_{\text{T}}-1},$ (10)
where $p$ denotes the index of the receive antenna in BS; $\lambda^{b}=c/f_{\mathrm{C}}^{b}$ is the wavelength with $c$ being the speed of light.
The sensing channel model on the $n$-th subcarrier during the $m$-th OFDM symbol time in the $b$-th CC is expressed as
$\mathbf{H}_{m,n}^{b}=\kappa_{S}^{b}e^{j2\pi\left(f_{\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{0}\right)}\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})\mathbf{a}_{\text{Tx}}^{\text{T}}(\theta_{\text{Tx}}),$ (11)
where $\mathbf{H}_{m,n}^{b}\in\mathbb{C}^{N_{\text{R}}\times N_{\text{T}}}$ and $\kappa_{S}^{b}=\sqrt{\frac{(\lambda^{b})^{2}}{(4\pi)^{3}r_{0}^{4}}}\beta_{S}^{b}$ is the attenuation between the target and BS with $\beta_{S}^{b}\sim\mathcal{CN}(0,\sigma_{\beta}^{2})$ being the target RCS [20]; $f_{\text{s}}^{b}=\frac{2f_{\mathrm{C}}^{b}v_{0}}{c}$ and $\tau_{0}=\frac{2r_{0}}{c}$ are the Doppler frequency shift and delay, respectively.
II-C2 Received sensing signal
At the receiver (Rx) of BS, the mixed echo signals on high and low-frequency bands are separated through matched filtering. Therefore, the received echo signal on the $n$-th subcarrier during the $m$-th OFDM symbol time is denoted by
$\mathbf{y}_{m,n}^{S}$ $=\mathbf{y}_{m,n}^{S,1}+\mathbf{y}_{m,n}^{S,2}$ (12)
$=\sum_{b=1}^{\mathcal{B}}\mathbf{H}_{m,n}^{b}\mathbf{x}_{m,n}^{b}+\mathbf{z}_{m,n}^{S},$
where $\mathbf{y}_{m,n}^{S}\in\mathbb{C}^{N_{\text{R}}\times 1}$ and $\mathbf{y}_{m,n}^{S,b}$ denotes the echo signal in $b$-th CC; $\mathbf{x}_{m,n}^{b}\in\mathbb{C}^{N_{\text{T}}\times 1}=\mathbf{X}_{n}^{b}\Upsilon_{m}$ is the transmit data with $\Upsilon_{m}=[0_{0},0_{1},\cdots,1_{m},\cdots,0_{M_{b}-1}]^{\text{T}}\in\mathbb{C}^{M_{b}\times 1}$; $\mathbf{z}_{m,n}^{S}\sim\mathcal{CN}(0,\sigma_{S}^{2})\in\mathbb{C}^{N_{\text{R}}\times 1}$ is an AWGN vector.
III A Novel CA-enabled MIMO-OFDM ISAC Signal Processing Method
Figure 2: CA-enabled MIMO-OFDM ISAC signal processing.
In this section, a novel CA-enabled MIMO-OFDM ISAC signal processing method is proposed, as shown in Fig. 2. Without loss of generality, we assume that the $1$-st CC is on the low-frequency band and the $2$-nd CC is on the high-frequency band. Through matched filtering at the Rx, the mixed echo signals on the two CCs are separated.
There are two challenges in sensing signal processing in the proposed ISAC system: 1) The initial phase misalignment of the echo signals on high and low-frequency bands due to attenuation and RCS; 2) The fusion of the sensing data on high and low-frequency bands with different physical-layer parameters, such as subcarrier spacing and symbol duration. Therefore, we consider both data-level and symbol-level fusion methods of the sensing data on high and low-frequency bands.
The data-level fusion method is as follows: After the separation of mixed echo signals on high and low-frequency bands, the joint angle-range-velocity estimation method proposed in [21] is applied to the echo signal on high-frequency band to yield the estimated range $r_{\text{data}}^{1}$ and velocity $v_{\text{data}}^{1}$ of target. Similarly, the estimated range $r_{\text{data}}^{2}$ and velocity $v_{\text{data}}^{2}$ of target are obtained by processing the echo signal on low-frequency band. Finally, the final estimation of range is $\hat{r}_{\text{data}}=(r_{\text{data}}^{1}+r_{\text{data}}^{2})/2$, while the final estimation of velocity is $\hat{v}_{\text{data}}=(v_{\text{data}}^{1}+v_{\text{data}}^{2})/2$. However, the sensing accuracy of data-level fusion method is low. Therefore, this paper proposes a two-stage symbol-level fusion method, including signal preprocessing stage and sensing information fusion stage, detailed in Section III-A.
III-A Signal Preprocessing Stage
Facing the above challenges, one of the purposes of sensing preprocessing is to align the initial phases of the echo signals on high and low-frequency bands.
III-A1 Angle estimation
With identical AoA and AoD for two CCs and high signal-to-noise ratio (SNR) of the echo signal on low-frequency band, we utilize the echo signal on low-frequency band to estimate AoA and AoD.
Firstly, we remove the transmit data $\mathbf{x}_{m,n}^{1}$ in $\mathbf{y}_{m,n}^{S,1}$ using known communication data, and the echo signal on low-frequency band by all antennas is
$\hat{\mathbf{H}}_{m,n}^{1}$ $=\mathbf{y}_{m,n}^{S,1}\left(\mathbf{x}_{m,n}^{1}\right)^{\text{H}}\left(\mathbf{x}_{m,n}^{1}\left(\mathbf{x}_{m,n}^{1}\right)^{\text{H}}\right)^{-1}$ (13)
$=\kappa_{S}^{b}e^{j2\pi\left(f_{\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{0}\right)}\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})\mathbf{a}_{\text{Tx}}^{\text{T}}(\theta_{\text{Tx}})\mathbf{I}+\mathbf{z}_{m,n}^{S,1},$
where $\hat{\mathbf{H}}_{m,n}^{1}\in\mathbb{C}^{N_{\text{R}}\times N_{\text{T}}}$ and $\mathbf{I}\in\mathbb{C}^{N_{\text{T}}\times N_{\text{T}}}$ is the identity matrix; $\mathbf{z}_{m,n}^{S,1}$ is an AWGN matrix.
Upon observing $\hat{\mathbf{H}}_{m,n}^{1}$, it is obvious that the AoA introduces a linear phase shift along the received antenna elements. The AoA estimation is obtained using the conventional multiple signal classification (MUSIC) method [22], with the following steps.
Step 1: Calculate the covariance matrix $\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}$ of signal $\hat{\mathbf{H}}_{m,n}^{1}$.
$\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}=\hat{\mathbf{H}}_{m,n}^{1}\left[\hat{\mathbf{H}}_{m,n}^{1}\right]^{\text{H}}.$ (14)
Step 2: $\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}$ is performed an eigenvalue decomposition (EVD) to obtain
$\text{eig}\left(\mathbf{R}_{\hat{\mathbf{H}}_{m,n}^{1}}\right)=\mathbf{U}_{\text{s}}\mathbf{\Lambda}_{\text{s}}\mathbf{U}_{\text{s}}^{\text{H}}+\mathbf{U}_{\text{n}}\mathbf{\Lambda}_{\text{n}}\mathbf{U}_{\text{n}}^{\text{H}},$ (15)
where $\mathbf{\Lambda}_{\text{s}}$ and $\mathbf{\Lambda}_{\text{n}}$ denote the diagonal matrix of signal and noise, respectively; $\mathbf{U}_{\text{s}}$ and $\mathbf{U}_{\text{n}}$ represent the subspaces of signal and noise, respectively. The [22] proves that the noise subspace is orthogonal to the signal subspace. Therefore, we traverse the signal subspace to obtain AoA estimation.
Step 3: Generate a MUSIC spatial spectral vector
$f_{\text{music}}(\theta)=\frac{1}{\mathbf{a}(\theta)^{\text{H}}\mathbf{U}_{\text{n}}\mathbf{U}_{\text{n}}^{\text{H}}\mathbf{a}(\theta)},$ (16)
where $\theta\in\left(0,\pi\right]$ and
$\mathbf{a}(\theta)=\left[e^{j2\pi p(\frac{d_{\text{r}}}{\lambda^{1}})\sin(\theta)}\right]^{\text{T}}|_{p=0,1,\cdots,N_{\text{R}}-1}.$ (17)
Step 4: Search the peak of $f_{\text{music}}$ to obtain the AoA estimation $\hat{\theta}_{\text{Rx}}$.
With the MUSIC method, the estimation result $\hat{\theta}_{\text{Rx}}$ of AoA is acquired for phase compensation.
III-A2 Phase compensation
Phase compensation is designed to cancel known communication data, AoA, and AoD information contained in the echo signals on high and low-frequency bands. The echo signal by all antennas on the $n$-th subcarrier during the $m$-th OFDM symbol time in the $b$-th CC is expressed as
$\mathbf{y}_{m,n}^{S,b}=\kappa_{S}^{b}e^{j2\pi\left(f_{\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{0}\right)}\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})\mathcal{X}_{m,n}^{b}+\mathbf{z}_{m,n}^{S,b},$ (18)
where $\mathbf{y}_{m,n}^{S,b}\in\mathbb{C}^{N_{\text{R}}\times 1}$ and $\mathcal{X}_{m,n}^{b}=\mathbf{a}_{\text{Tx}}^{\text{T}}(\theta_{\text{Tx}})\mathbf{x}_{m,n}^{b}\in\mathbb{C}$ is a transmit factor.
With the known transmit data and AoD estimation, we can compensate for $\mathcal{X}_{m,n}^{b}$ by multiplying the compensation factor $\widetilde{\mathcal{X}}_{m,n}^{b}=\left(\mathbf{a}_{\text{Tx}}^{\text{T}}(\hat{\theta}_{\text{Tx}})\mathbf{x}_{m,n}^{b}\right)^{*}$ on both sides of (18), denoted by
$\mathbf{y}_{m,n}^{S,b}\widetilde{\mathcal{X}}_{m,n}^{b}=$ (19)
$\hat{\mathcal{X}}_{m,n}^{b}\kappa_{S}^{b}e^{j2\pi\left(f_{\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{0}\right)}\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})+\mathbf{z}_{m,n}^{S,b}\widetilde{\mathcal{X}}_{m,n}^{b},$
where $\hat{\mathcal{X}}_{m,n}^{b}=\mathcal{X}_{m,n}^{b}\widetilde{\mathcal{X}}_{m,n}^{b}\in\mathbb{R}$. Then, with the AoA estimation, $\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})$ in (19) is compensated, and $\mathbf{r}_{m,n}^{S,b}$ is expressed as
$\mathbf{r}_{m,n}^{S,b}=\mathbf{h}_{m,n}^{S,b}+\mathbf{n}_{m,n}^{S,b},$ (20)
where
$\mathbf{r}_{m,n}^{S,b}\in\mathbb{C}^{N_{\text{R}}\times 1}=\mathbf{y}_{m,n}^{S,b}\widetilde{\mathcal{X}}_{m,n}^{b}\circ\mathbf{a}_{\text{Rx}}^{*}(\theta_{\text{Rx}}),$ (21)
$\mathbf{n}_{m,n}^{S,b}=\mathbf{z}_{m,n}^{S,b}\widetilde{\mathcal{X}}_{m,n}^{b}\circ\mathbf{a}_{\text{Rx}}^{*}(\theta_{\text{Rx}}),$ (22)
$\mathbf{h}_{m,n}^{S,b}=\hat{\mathcal{X}}_{m,n}^{b}\kappa_{S}^{b}e^{j2\pi\left(f_{\text{s}}^{b}mT^{b}-n\Delta f^{b}\tau_{0}\right)}\mathbf{a}_{\text{Rx}}(\theta_{\text{Rx}})\circ\mathbf{a}_{\text{Rx}}^{*}(\theta_{\text{Rx}}).$ (23)
III-A3 Space-domain diversity
Upon observing Eq. (23), the echo signals on each antenna are identical and can be coherently accumulated to improve SNR. The echo signal after coherent accumulation on the $n$-th subcarrier during the $m$-th OFDM symbol time in the $b$-th CC is expressed as
$\mathbf{d}_{m,n}^{S,b}=\frac{1}{N_{\text{R}}}\sum_{p=0}^{N_{\text{R}}-1}\mathbf{r}_{m,n}^{S,b}(p),$ (24)
where $\mathbf{r}_{m,n}^{S,b}(p)$ is the $p$-th element of $\mathbf{r}_{m,n}^{S,b}$. According to (24), the echo signal on $N_{b}$ subcarriers during $M_{b}$ OFDM symbol times in the $b$-th CC is written in matrix form as
$\mathbf{D}^{S,b}=\mathbf{S}^{S,b}+\mathbf{Z}^{S,b},$ (25)
where $\mathbf{Z}^{S,b}$ is an AWGN matrix and $\mathbf{S}^{S,b}\in\mathbb{C}^{N_{b}\times M_{b}}$ denotes a delay-Doppler information matrix, expressed in (26).
$\centering\mathbf{S}^{S,b}=\hat{\mathcal{X}}_{m,n}^{b}\kappa_{S}^{b}\left[\begin{array}[]{cccc}1&e^{j2\pi f_{\text{s}}^{b}T^{b}}&\cdots&e^{j2\pi(M_{b}-1)f_{\text{s}}^{b}T^{b}}\\ e^{-j2\pi\Delta f^{b}\tau_{0}}&e^{-j2\pi\Delta f^{b}\tau_{0}}e^{j2\pi f_{\text{s}}^{b}T^{b}}&\cdots&e^{-j2\pi\Delta f^{b}\tau_{0}}e^{j2\pi(M_{b}-1)f_{\text{s}}^{b}T^{b}}\\ \vdots&\vdots&\ddots&\vdots\\ e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{0}}&e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{0}}e^{j2\pi f_{\text{s}}^{b}T^{b}}&\cdots&e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{0}}e^{j2\pi(M_{b}-1)f_{\text{s}}^{b}T^{b}}\\ \end{array}\right].\@add@centering$ (26)
III-A4 CCC method
Observing (26), the initial phases of $\mathbf{S}^{S,b}$ on high and low-frequency bands are not aligned and there exists a different complex number $\hat{\mathcal{X}}_{m,n}^{b}\kappa_{S}^{b}$. To this end, a CCC method [23] is performed to eliminate the complex number $\hat{\mathcal{X}}_{m,n}^{b}\kappa_{S}^{b}$. Considering the noise, the procedure for applying the CCC method to $\mathbf{D}^{S,b}$ is presented in the following steps.
Step 1: $\mathbf{D}^{S,b}$ is divided into $N_{b}$ row vectors, denoted by
$\mathbf{D}^{S,b}=\left[\left(\mathbf{D}_{0}^{S,b}\right)^{\text{T}},\left(\mathbf{D}_{1}^{S,b}\right)^{\text{T}},\cdots,\left(\mathbf{D}_{N_{b}-1}^{S,b}\right)^{\text{T}}\right]^{\text{T}},$ (27)
where $\mathbf{D}_{n}^{S,b}\in\mathbb{C}^{1\times M_{b}}$ represents the $n$-th row vector of $\mathbf{D}^{S,b}$.
Step 2: A delay feature vector $\mathbf{R}_{S}^{b}\in\mathbb{C}^{N_{b}\times 1}$ is obtained by accumulating the results of conjugate multiplication between the row vectors, denoted by (28), where $\mathbf{z}^{b}$ is an AWGN vector.
$\mathbf{R}_{S}^{b}$ $=\frac{1}{M_{b}}\left[\frac{1}{N_{b}}\sum_{a=0}^{N_{b}-1}\mathbf{D}_{a}^{S,b}\left(\mathbf{D}_{a}^{S,b}\right)^{\text{H}},\frac{1}{N_{b}-1}\sum_{a=0}^{N_{b}-1-1}\mathbf{D}_{a+1}^{S,b}\left(\mathbf{D}_{a}^{S,b}\right)^{\text{H}},\cdots,\frac{1}{N_{b}-n}\sum_{a=0}^{N_{b}-n-1}\mathbf{D}_{a+n}^{S,b}\left(\mathbf{D}_{a}^{S,b}\right)^{\text{H}},\cdots,\mathbf{D}_{N_{b}-1}^{S,b}\left(\mathbf{D}_{0}^{S,b}\right)^{\text{H}}\right]^{\text{T}}$ (28)
$=\left[1,e^{-j2\pi\Delta f^{b}\tau_{0}},\cdots,e^{-j2\pi n\Delta f^{b}\tau_{0}},\cdots,e^{-j2\pi(N_{b}-1)\Delta f^{b}\tau_{0}}\right]^{\text{T}}+\mathbf{z}^{b}.$
Step 3: $\mathbf{D}^{S,b}$ is divided into $M_{b}$ column vectors, expressed as
$\mathbf{D}^{S,b}=\left[\mathbf{C}_{0}^{S,b},\mathbf{C}_{1}^{S,b},\cdots,\mathbf{C}_{M_{b}-1}^{S,b}\right],$ (29)
where $\mathbf{C}_{m}^{S,b}\in\mathbb{C}^{N_{b}\times 1}$ denotes the $m$-th column vector of $\mathbf{D}^{S,b}$.
Step 4: A Doppler feature vector $\mathbf{V}_{S}^{b}\in\mathbb{C}^{1\times M_{b}}$ is obtained by accumulating the result of conjugate multiplication between the column vectors, expressed as (30).
$\mathbf{V}_{S}^{b}$ $=\frac{1}{N_{b}}\left[\frac{1}{M_{b}}\sum_{a=0}^{M_{b}-1}\left(\mathbf{C}_{a}^{S,b}\right)^{\text{H}}\mathbf{C}_{a}^{S,b},\frac{1}{M_{b}-1}\sum_{a=0}^{M_{b}-1-1}\left(\mathbf{C}_{a}^{S,b}\right)^{\text{H}}\mathbf{C}_{a+1}^{S,b},\cdots,\frac{1}{M_{b}-n}\sum_{a=0}^{M_{b}-n-1}\left(\mathbf{C}_{a}^{S,b}\right)^{\text{H}}\mathbf{C}_{a+n}^{S,b},\cdots,\left(\mathbf{C}_{0}^{S,b}\right)^{\text{H}}\mathbf{C}_{N_{b}-1}^{S,b}\right]$ (30)
$=\left[1,e^{j2\pi\left(\frac{2f_{\mathrm{C}}^{b}v_{0}}{c}\right)T^{b}},\cdots,e^{j2\pi m\left(\frac{2f_{\mathrm{C}}^{b}v_{0}}{c}\right)T^{b}},\cdots,e^{j2\pi(M_{b}-1)\left(\frac{2f_{\mathrm{C}}^{b}v_{0}}{c}\right)T^{b}}\right]+\mathbf{\hat{z}}^{b}.$
After the above steps, the four feature vectors are obtained, namely $\mathbf{R}_{S}^{1}$, $\mathbf{V}_{S}^{1}$, $\mathbf{R}_{S}^{2}$, and $\mathbf{V}_{S}^{2}$.
III-B Sensing Information Fusion Stage
In the sensing information fusion stage, we fuse the feature vectors obtained in the signal preprocessing stage by the symbol-level fusion and estimate the parameters of target, which include the fusion algorithm of delay feature vectors and the fusion algorithm of Doppler feature vectors. Without loss of generality, the following remarks need to be specified.
[…]
  • Upon observing Fig. 2, the two CCs use discrete Add CP modules, allowing for the adjustment of CP lengths. It is imperative to ensure that the maximum delay does not surpass the length of the shortest CP [24, 7].
III-B1 Fusion algorithm of delay feature vectors
Without considering the noise, observing (28), the difference between $\mathbf{R}_{S}^{1}$ and $\mathbf{R}_{S}^{2}$ is the subcarrier spacing, which makes the symbol-level fusion difficult. According to $\frac{\Delta f^{2}}{\Delta f^{1}}=Q$, $\mathbf{R}_{S}^{1}$ and $\mathbf{R}_{S}^{2}$ are rewritten as
$\mathbf{R}_{S}^{1}$ $=\left[e^{-j2\pi n\Delta f^{1}\tau_{0}}\right]^{\text{T}}|_{n=0,1,\cdots,N_{1}-1},$ (31)
$\mathbf{R}_{S}^{2}$ $=\left[e^{-j2\pi nQ\Delta f^{1}\tau_{0}}\right]^{\text{T}}|_{n=0,1,\cdots,N_{2}-1}.$ (32)
(32) is further rewritten as
$\hat{\mathbf{R}}_{S}^{2}=\left[e^{-j2\pi n^{\prime\prime}\Delta f^{1}\tau_{0}}\right]^{\text{T}}\in\mathbb{C}^{N_{2}\times 1}|_{n^{\prime\prime}=0,Q,\cdots,Q(N_{2}-1)}.$ (33)
Then, the subcarrier spacings in (31) and (33) are identical. The fusion process of delay feature vectors is discussed below in two cases.
Case 1 $(N_{1}-1<Q(N_{2}-1))$: Initialize a zero-element vector $\mathbf{P}_{S}^{1}\in\mathbb{C}^{QN_{2}\times 1}$ and an index value $\xi_{1}\in\{0,1,\cdots,QN_{2}-1\}$. Then, we traverse the element of $\mathbf{P}_{S}^{1}$ with $\xi_{1}$. When $\xi_{1}$ equals $n$, the element $\mathbf{R}_{S}^{1}(n)$ is assigned to $\mathbf{P}_{S}^{1}(\xi_{1})$. When $\xi_{1}$ equals $n^{\prime\prime}$, the element $\hat{\mathbf{R}}_{S}^{2}(n^{\prime\prime})$ is assigned to $\mathbf{P}_{S}^{1}(\xi_{1})$. Given that the zero elements in $\mathbf{P}_{S}^{1}$ deteriorate the Fourier sidelobes, which can be mitigated by recovering the missing sensing data with the CS-based sensing method proposed in [11]. The final processed vector is denoted by $\hat{\mathbf{P}}_{S}^{1}$.
Case 2 $\left(N_{1}-1\geq Q(N_{2}-1)\right)$: Initialize a zero-element vector $\mathbf{P}_{S}^{2}\in\mathbb{C}^{N_{1}\times 1}$ and an index value $\xi_{2}\in\{0,1,\cdots,N_{1}-1\}$. Then, we traverse the element of $\mathbf{P}_{S}^{2}$ with $\xi_{2}$. When $\xi_{2}$ equals $n$, the element $\mathbf{R}_{S}^{1}(n)$ is assigned to $\mathbf{P}_{S}^{2}(\xi_{2})$. When $\xi_{2}$ equals $n^{\prime\prime}$, the element $\hat{\mathbf{R}}_{S}^{2}(n^{\prime\prime})$ is assigned to $\mathbf{P}_{S}^{2}(\xi_{2})$.
Taking the $\hat{\mathbf{P}}_{S}^{1}$ in Case 1 as an example, an improved IDFT method is used to achieve a high-accuracy sensing. The specific procedures are as follows, while the Case 2 has the similar procedures.
Step 1: Obtain a searching interval $\left[R_{\text{min}},R_{\text{max}}\right]$ based on the sensing demands in the application scenario of ISAC [25].
Step 2: Generate a distance searching vector $\vec{\mathbf{A}}$ by gridding the searching interval with a size of grid being $\Delta R=\frac{R_{\text{max}}-R_{\text{min}}}{J}$, where $J$ is the number of girds and $\vec{\mathbf{A}}\in\mathbb{C}^{J\times 1}$ is denoted by
$\vec{\mathbf{A}}=\left[R_{1},R_{2},\cdots,R_{\eta},\cdots,R_{J}\right]^{\text{T}},$ (34)
with $\eta\in\{1,2,\cdots,J\}$ denoting the index of grids.
Step 3: According to the size of $\hat{\mathbf{P}}_{S}^{1}$ and the expression form of delay feature vectors in (31), the $\vec{\mathbf{A}}$ is transformed to a searching matrix $\mathbf{A}_{1}\in\mathbb{C}^{J\times QN_{2}}$, which is expressed as
$\mathbf{A}_{1}=\left[\begin{array}[]{cccc}1&e^{j2\pi\Delta f^{1}\frac{2R_{1}}{c}}&\cdots&e^{j2\pi(QN_{2}-1)\Delta f^{1}\frac{2R_{1}}{c}}\\ 1&e^{j2\pi\Delta f^{1}\frac{2R_{2}}{c}}&\cdots&e^{j2\pi(QN_{2}-1)\Delta f^{1}\frac{2R_{2}}{c}}\\ \vdots&\vdots&\ddots&\vdots\\ 1&e^{j2\pi\Delta f^{1}\frac{2R_{J}}{c}}&\cdots&e^{j2\pi(QN_{2}-1)\Delta f^{1}\frac{2R_{J}}{c}}\\ \end{array}\right].$ (35)
Step 4: Obtain and search the delay profile $\vec{\mathbf{P}}_{1}=\mathbf{A}_{1}\hat{\mathbf{P}}_{S}^{1}$. The peak index of $\vec{\mathbf{P}}_{1}$ is denoted by $\hat{\eta}_{1}$.
Step 5: The range estimation of target is $\hat{r}_{0}^{1}=\vec{\mathbf{A}}(\hat{\eta}_{1})$.
The fusion algorithm of delay feature vectors is shown in Algorithm 1, which is explained intuitively in Fig. 3.
Figure 3: Fusion algorithm of delay feature vectors when Q = 2.
Algorithm 1: Fusion algorithm of Delay feature vectors
Input:
Delay feature vectors $\mathbf{R}_{S}^{1}$ and $\mathbf{R}_{S}^{2}$;
The number of subcarriers $N_{1}$ and $N_{2}$;
The subcarriers spacing $\Delta f^{1}$ and $\Delta f^{2}$.
Output: The range estimations of target $\hat{r}_{0}^{1}$ and $\hat{r}_{0}^{2}$.
1: if $N_{1}-1\geq Q(N_{2}-1)$ do
2:     Assume a vector $\mathbf{P}_{S}^{2}\in\mathbb{C}^{N_{1}\times 1}=\mathbf{R}_{S}^{1}$,
    and an index value $\mu_{2}\in\mathcal{N}_{2}=\{0,1,\cdots,N_{2}-1\}$;
3:     for $\mathbf{P}_{S}^{2}$ $\mu_{2}$ in $\mathcal{N}_{2}$ do
4:        The element value of $\mathbf{R}_{S}^{2}(\mu_{2})$ is added to $\mathbf{P}_{S}^{2}(Q\mu_{2})$;
5:        $\mathbf{P}_{S}^{2}(Q\mu_{2})$ = $\mathbf{P}_{S}^{2}(Q\mu_{2})/2$;
6:     end for
7: else if $N_{1}-1<Q(N_{2}-1)$ do
8:     Initialize a zero-element vector $\mathbf{P}_{S}^{1}\in\mathbb{C}^{QN_{2}\times 1}$,
    and an index value $\mu_{1}\in\mathcal{N}_{1}=\{0,1,\cdots,N_{1}-1\}$;
9:     for $\mathbf{P}_{S}^{1}$ $\mu_{1}$ in $\mathcal{N}_{1}$ do
10:        The element value of $\mathbf{R}_{S}^{1}(\mu_{1})$ is assigned to $\mathbf{P}_{S}^{1}(\mu_{1})$;
11:     end for
12:     Initialize an index value $\mu_{3}\in\mathcal{N}_{3}=\{0,1,\cdots,N_{2}-1\}$;
13:     for $\mathbf{P}_{S}^{1}$ $\mu_{3}$ in $\mathcal{N}_{3}$ do
14:        The element value of $\mathbf{R}_{S}^{2}(\mu_{3})$ is added to $\mathbf{P}_{S}^{1}(Q\mu_{3})$;
15:        if $Q\mu_{3}\leq N_{1}-1$ do
16:           $\mathbf{P}_{S}^{1}(Q\mu_{3})$ = $\mathbf{P}_{S}^{1}(Q\mu_{3})/2$;
17:        end if
18:     end for
19:     Recover the missing elements of $\mathbf{P}_{S}^{1}$ by
    CS-based method proposed in [11] to obtain $\hat{\mathbf{P}}_{S}^{1}$;
20: end if
21: The results of $\mathbf{P}_{S}^{1}$ and $\mathbf{P}_{S}^{2}$ undergoing improved IDFT are
searched for peaks to obtain the peak indices $\hat{\eta}^{1}$ and $\hat{\eta}^{2}$;
22: Substitute $\hat{\eta}^{1}$ and $\hat{\eta}^{2}$ into (34) to obtain the range
estimations of target $\hat{r}_{0}^{1}$ and $\hat{r}_{0}^{2}$.
III-B2 Fusion algorithm of Doppler feature vectors
Without considering the noise, observing (30), the different parameters between $\mathbf{V}_{S}^{1}$ and $\mathbf{V}_{S}^{2}$ are carrier frequency and total symbol duration. (30) is rewritten as
$\mathbf{V}_{S}^{b}=\left[e^{j2\pi m\frac{2v_{0}}{c}\left(f_{\mathrm{C}}^{b}T^{b}\right)}\right]|_{m=0,1,\cdots,M_{b}-1},$ (36)
where $f_{\mathrm{C}}^{b}T^{b}=f_{\mathrm{C}}^{b}\left(\frac{1}{\Delta f^{b}}+T_{\text{s}}^{b}\right)$. If $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$, $\mathbf{V}_{S}^{1}$ and $\mathbf{V}_{S}^{2}$ can be coherently accumulated to obtain a SNR gain.
Fortunately, as shown in Section III-B, the unique framework of the proposed CA-enabled MIMO-OFDM ISAC system supports adjustable durations of CPs in the two CCs. Meanwhile, Theorem 1 proves that $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$.
Theorem 1
According to $\left\lfloor\frac{f_{\mathrm{C}}^{2}}{f_{\mathrm{C}}^{1}}\right\rfloor=Q$, we assume that $\frac{f_{\mathrm{C}}^{2}}{f_{\mathrm{C}}^{1}}=Q+\rho(\rho<1)$. $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$ can be satisfied when
$N_{\text{s}}^{1}=\left(1+\frac{\rho}{Q}\right)\frac{N_{1}}{N_{2}}N_{\text{s}}^{2}+\frac{N_{1}\rho}{Q}.$ (37)
Proof:
Based on $\frac{f_{\mathrm{C}}^{2}}{f_{\mathrm{C}}^{1}}=\frac{\Delta f^{2}}{\Delta f^{1}}+\rho$, the relationship between $T_{\text{s}}^{1}$ and $T_{\text{s}}^{2}$ is expressed as
$f_{\mathrm{C}}^{1}\left(\frac{1}{\Delta f^{1}}+T_{\text{s}}^{1}\right)$ $=f_{\mathrm{C}}^{2}\left(\frac{1}{\Delta f^{2}}+T_{\text{s}}^{2}\right),$ (38)
$\Rightarrow\left(\frac{f_{\mathrm{C}}^{1}}{\Delta f^{1}}+f_{\mathrm{C}}^{1}T_{\text{s}}^{1}\right)$ $=f_{\mathrm{C}}^{1}\left(\frac{\Delta f^{2}}{\Delta f^{1}}+\rho\right)\left(\frac{1}{\Delta f^{2}}+T_{\text{s}}^{2}\right),$
$\Rightarrow T_{\text{s}}^{1}$ $=(Q+\rho)T_{\text{s}}^{2}+\frac{\rho}{\Delta f^{2}}.$
Then, according to $T_{\text{s}}^{b}=\frac{N_{\text{s}}^{b}T_{\text{ofdm}}^{b}}{N_{b}}$, the relationship between $N_{\text{s}}^{1}$ and $N_{\text{s}}^{2}$ can be expressed as
$\frac{N_{\text{s}}^{1}T_{\text{ofdm}}^{1}}{N_{1}}$ $=(Q+\rho)\frac{N_{\text{s}}^{2}T_{\text{ofdm}}^{2}}{N_{2}}+\frac{\rho}{\Delta f^{2}},$ (39)
$\Rightarrow N_{\text{s}}^{1}$ $=(\frac{Q+\rho}{Q})\frac{N_{1}}{N_{2}}N_{\text{s}}^{2}+\frac{N_{1}\rho}{Q}.$
The fusion process of Doppler feature vectors is revealed in the following three cases.
Case 1 $(M_{1}>M_{2})$: Initialize a zero-element vector $\mathbf{E}_{S}^{1}\in\mathbb{C}^{1\times M_{1}}$ and complement $\mathbf{V}_{S}^{2}$ with zero to get the matrix $\hat{\mathbf{V}}_{S}^{2}\in\mathbb{C}^{1\times M_{1}}=[\mathbf{V}_{S}^{2},0_{M_{2}},\cdots,0_{M_{1}-1}]$. Then, the processed vector $\mathbf{E}_{S}^{1}=(\mathbf{V}_{S}^{1}+\hat{\mathbf{V}}_{S}^{2})/2$ is obtained.
Case 2 $(M_{1}<M_{2})$: Initialize a zero-element vector $\mathbf{E}_{S}^{2}\in\mathbb{C}^{1\times M_{2}}$ and complement $\mathbf{V}_{S}^{1}$ with zero to get the matrix $\hat{\mathbf{V}}_{S}^{1}\in\mathbb{C}^{1\times M_{2}}=[\mathbf{V}_{S}^{1},0_{M_{1}},\cdots,0_{M_{2}-1}]$. Then, the processed vector $\mathbf{E}_{S}^{2}=(\hat{\mathbf{V}}_{S}^{1}+\mathbf{V}_{S}^{2})/2$ is obtained.
Case 3 $(M_{1}=M_{2}=M)$: Initialize a zero-element vector $\mathbf{E}_{S}^{3}\in\mathbb{C}^{1\times M}$ and obtain a vector $\mathbf{E}_{S}^{3}=(\mathbf{V}_{S}^{1}+\mathbf{V}_{S}^{2})/2$.
Taking the $\mathbf{E}_{S}^{1}$ in Case 1 as an example, an improved discrete Fourier transform (DFT) method is applied to achieve high-accuracy sensing. The specific procedures are as follows, while the Case 2 and Case 3 have the similar procedures.
Step 1: Obtain a searching interval $\left[V_{\text{min}},V_{\text{max}}\right]$ based on the sensing demands in the application scenario of ISAC [25].
Step 2: Generate a velocity searching vector $\vec{\mathbf{B}}$ by gridding the searching interval with a size of grid being $\Delta V=\frac{V_{\text{max}}-V_{\text{min}}}{G}$, where $G$ is the number of grids and $\vec{\mathbf{B}}\in\mathbb{C}^{1\times G}$ is expressed as
$\vec{\mathbf{B}}=\left[V_{1},V_{2},\cdots,V_{g},\cdots,V_{G}\right],$ (40)
with $g\in\{1,2,\cdots,G\}$ representing the index of grids.
Step 3: According to the size of $\mathbf{E}_{S}^{1}$ and the expression form of Doppler feature vectors in (30), $\vec{\mathbf{B}}$ is transformed to a searching matrix $\mathbf{B}_{1}\in\mathbb{C}^{M_{1}\times G}$, which is expressed as
$\mathbf{B}_{1}=\left[\begin{array}[]{cccc}1&e^{-j2\pi f_{\mathrm{C}}^{1}T^{1}\frac{2V_{1}}{c}}&\cdots&e^{-j2\pi(M_{1}-1)f_{\mathrm{C}}^{1}T^{1}\frac{2V_{1}}{c}}\\ 1&e^{-j2\pi f_{\mathrm{C}}^{1}T^{1}\frac{2V_{2}}{c}}&\cdots&e^{-j2\pi(M_{1}-1)f_{\mathrm{C}}^{1}T^{1}\frac{2V_{2}}{c}}\\ \vdots&\vdots&\ddots&\vdots\\ 1&e^{-j2\pi f_{\mathrm{C}}^{1}T^{1}\frac{2V_{G}}{c}}&\cdots&e^{-j2\pi(M_{1}-1)f_{\mathrm{C}}^{1}T^{1}\frac{2V_{G}}{c}}\\ \end{array}\right]^{\text{T}}.$ (41)
Step 4: Obtain and search the Doppler profile $\vec{\mathbf{E}}_{1}=\mathbf{E}_{S}^{1}\mathbf{B}_{1}$. The peak index of $\vec{\mathbf{E}}_{1}$ is denoted by $\hat{g}_{1}$.
Step 5: The velocity estimation of target is $\hat{v}_{0}^{1}=\vec{\mathbf{B}}(\hat{g}_{1})$.
The fusion algorithm of Doppler feature vectors is shown in Algorithm 2, which is explained intuitively in Fig. 4.
Algorithm 2 : Fusion algorithm of Doppler feature vectors
Input:
Doppler feature vectors $\mathbf{V}_{S}^{1}$ and $\mathbf{V}_{S}^{2}$;
The number of OFDM symbols $M_{1}$ and $M_{2}$;
The carrier frequencies $f_{\mathrm{C}}^{1}$ and $f_{\mathrm{C}}^{2}$;
The total OFDM duration $T^{1}$ and $T^{2}$.
Output: The velocity estimations of target $\hat{v}_{0}^{1}$, $\hat{v}_{0}^{2}$, and $\hat{v}_{0}^{3}$.
1: if $M_{1}>M_{2}$ do
2:     Assume a vector $\mathbf{E}_{S}^{1}\in\mathbb{C}^{1\times M_{1}}=\mathbf{V}_{S}^{1}$,
    and an index value $\omega_{1}\in\mathcal{M}_{1}=\{0,1,\cdots,M_{2}-1\}$;
3:     for $\mathbf{E}_{S}^{1}$ $\omega_{1}$ in $\mathcal{M}_{1}$ do
4:        The element value of $\mathbf{V}_{S}^{2}(\omega_{1})$ is added to $\mathbf{E}_{S}^{1}(\omega_{1})$;
5:        $\mathbf{E}_{S}^{1}(\omega_{1})=\mathbf{E}_{S}^{1}(\omega_{1})/2$;
6:     end for
7: else if $M_{1}<M_{2}$ do
8:     Assume a vector $\mathbf{E}_{S}^{2}\in\mathbb{C}^{1\times M_{2}}=\mathbf{V}_{S}^{2}$,
    and an index value $\omega_{2}\in\mathcal{M}_{2}=\{0,1,\cdots,M_{1}-1\}$;
9:     for $\mathbf{E}_{S}^{2}$ $\omega_{2}$ in $\mathcal{M}_{2}$ do
10:        The element value of $\mathbf{V}_{S}^{1}(\omega_{2})$ is added to $\mathbf{E}_{S}^{2}(\omega_{2})$;
11:        $\mathbf{E}_{S}^{2}(\omega_{2})=\mathbf{E}_{S}^{2}(\omega_{2})/2$;
12:     end for
13: else if $M_{1}=M_{2}=M$ do
14:     Initialize a zero-elements vector $\mathbf{E}_{S}^{3}\in\mathbb{C}^{1\times M}$,
    and an index value $\omega_{3}\in\mathcal{M}_{3}=\{0,1,\cdots,M-1\}$;
15:     for $\mathbf{E}_{S}^{3}$ $\omega_{3}$ in $\mathcal{M}_{3}$ do
16:        $\mathbf{E}_{S}^{3}(\omega_{3})=\left[\mathbf{V}_{S}^{1}(\omega_{3})+\mathbf{V}_{S}^{2}(\omega_{3})\right]/2$;
17:     end for
18: end if
19: The results of $\mathbf{E}_{S}^{1}$, $\mathbf{E}_{S}^{2}$, and $\mathbf{E}_{S}^{3}$ undergoing improved DFT are
searched for peaks to obtain the peak indices $\hat{g}_{1}$, $\hat{g}_{2}$, and $\hat{g}_{3}$;
20: Substitute $\hat{g}_{1}$, $\hat{g}_{2}$, and $\hat{g}_{3}$ into (40) to obtain the velocity
estimations of target $\hat{v}_{0}^{1}$, $\hat{v}_{0}^{2}$, and $\hat{v}_{0}^{3}$.
Figure 4: Fusion algorithm of Doppler feature vectors.
IV Performance Analysis
In this section, communication MI is used to characterize the channel capacity of the proposed ISAC system, and CRLB is used as a performance metric of sensing for the proposed ISAC system [26].
IV-A Communication Performance Analysis
In this section, the closed-form communication MI of the proposed CA-enabled MIMO-OFDM ISAC system is derived with the assumption of known communication CSI at both the BS and UE.
IV-A1 Communication MI of MIMO-OFDM ISAC system
The communication MI for MIMO-OFDM ISAC system is provided in Lemma 1.
Lemma 1
According to the signal model in (42), the communication MI of the MIMO-OFDM ISAC system is obtained in (43), with $N_{x}$, $N_{c}$, and $N_{r}$ denoting the number of OFDM symbols, subcarriers, and received antennas, respectively [27].
$\mathbf{Y}_{\text{com}}=\mathbf{X}\mathbf{H}+\mathbf{W}_{\text{com}},$ (42)
$I(\mathbf{X};\mathbf{Y}_{\text{com}}{|\mathbf{H})}$ (43)
$=N_{x}\mathrm{log}_{2}\left[\left(\sigma_{\text{Z}}^{2}\right)^{-N_{r}N_{c}}\det(\mathbf{H}^{\text{H}}\frac{E\{\mathbf{X}^{\text{H}}\mathbf{X}\}}{N_{x}}\mathbf{H}+\sigma_{\text{Z}}^{2}\mathbf{I}_{N_{r}N_{c}})\right],$
where $\mathbf{W}_{\text{com}}\sim\mathcal{CN}\left(0,\sigma_{\text{Z}}^{2}\right)$.
IV-A2 Communication MI of the proposed CA-enabled MIMO-OFDM ISAC system
According to [27] and (8), the communication MI of the proposed CA-enabled MIMO-OFDM ISAC system is defined as
$I\left(\mathbf{X}_{\mathrm{C}}^{1},\mathbf{X}_{\mathrm{C}}^{2};\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{1}\right).$ (44)
Since the frequency bands between the two CCs are far separated and transmit data independently, $\mathbf{X}_{\mathrm{C}}^{1}$ and $\mathbf{X}_{\mathrm{C}}^{2}$ are independent, as well as $\mathbf{H}_{\mathrm{C}}^{1}$ and $\mathbf{H}_{\mathrm{C}}^{2}$. Therefore, (44) is rewritten as
$I\left(\mathbf{X}_{\mathrm{C}}^{1},\mathbf{X}_{\mathrm{C}}^{2};\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)$ (45)
$=I\left(\mathbf{X}_{\mathrm{C}}^{1};\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)+I\left(\mathbf{X}_{\mathrm{C}}^{2};\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{X}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)$
$=h\left(\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)-h\left(\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{X}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)$
$\quad+h\left(\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{X}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)-h\left(\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{X}_{\mathrm{C}}^{1},\mathbf{X}_{\mathrm{C}}^{2},\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)$
$=h{\left(\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)}-h{\left(\mathbf{W}_{u}^{\mathrm{C}}\right)}.$
Theorem 2
According to (8) and (45), the communication MI of the proposed CA-enabled MIMO-OFDM ISAC system is expressed as
$I\left(\mathbf{X}_{\mathrm{C}}^{1},\mathbf{X}_{\mathrm{C}}^{2};\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)=$ (46)
$M\log_{2}\left[\left(\sigma_{\mathrm{C}}^{2}\right)^{-N_{\text{U}}N}\det\left(\sum_{b=1}^{\mathcal{B}}\left[\left(\mathbf{H}^{b}\right)^{\text{H}}E\left[\frac{\left(\mathbf{X}_{\mathrm{C}}^{b}\right)^{\text{H}}\mathbf{X}_{\mathrm{C}}^{b}}{M}\right]\mathbf{H}^{b}\right]+\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}N}\right)\right].$
Proof:
For ease of derivation, we consider the assumptions of $N_{1}=N_{2}=N$ and $M_{1}=M_{2}=M$. Since $\mathbf{X}_{\mathrm{C}}^{b}$ is a block diagonal array, the probability density function (PDF) of $\mathbf{Y}_{u}^{\mathrm{C}}$ under known CSI is the product of the PDFs on the $N$ subcarriers, denoted by (47), where $\mathbf{y}_{u}^{\mathrm{C},j}$ is the $j$-th row or column vector of $\mathbf{Y}_{u}^{\mathrm{C}}$, and $\mathbf{I}_{N_{\text{U}}}$ represents a $N_{\text{U}}\times N_{\text{U}}$ unit matrix.
$p\left(\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)=\prod_{j=1}^{NM}p\left(\mathbf{y}_{u}^{\mathrm{C},j}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)$ (47)
$=\prod_{j=1}^{NM}\frac{1}{\pi^{N_{\text{U}}}\text{det}\left[\sum\limits_{b=1}^{\mathcal{B}}\left[\left(\mathbf{H}_{\mathrm{C}}^{b}\right)^{\text{H}}E\left\{\left(\mathbf{X}_{\mathrm{C}}^{b,j}\right)^{\text{H}}\mathbf{X}_{\mathrm{C}}^{b,j}\right\}\mathbf{H}_{\mathrm{C}}^{b}\right]+\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}}\right]}\exp\left\{-\left[\left(\sum\limits_{b=1}^{\mathcal{B}}\left[\left(\mathbf{H}_{\mathrm{C}}^{b}\right)^{\text{H}}E\left\{\left(\mathbf{X}_{\mathrm{C}}^{b,j}\right)^{\text{H}}\mathbf{X}_{\mathrm{C}}^{b,j}\right\}\mathbf{H}_{\mathrm{C}}^{b}\right]+\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}}\right)^{-1}\mathbf{y}_{u}^{\mathrm{C},j}\left(\mathbf{y}_{u}^{\mathrm{C},j}\right)^{\text{H}}\right]\right\}$
$=\prod_{j=1}^{N}\frac{1}{\pi^{MN_{\text{U}}}\det^{M}\left[\sum\limits_{b=1}^{\mathcal{B}}\left[\left(\tilde{\mathbf{H}}_{u,m}^{n}\right)^{\text{H}}E\left\{\left(\mathbf{X}_{n}^{b,j}\right)^{\text{H}}\mathbf{X}_{n}^{b,j}\right\}\tilde{\mathbf{H}}_{a,m}^{n}\right]+\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}}\right]}$
×$\quad\times\exp\left\{-\text{tr}\left[\sum\limits_{j=1}^{M}\left(\sum\limits_{b=1}^{\mathcal{B}}\left[\left(\tilde{\mathbf{H}}_{u,m}^{n}\right)^{\text{H}}E\left\{\left(\mathbf{X}_{n}^{b,j}\right)^{\text{H}}\mathbf{X}_{n}^{b,j}\right\}\tilde{\mathbf{H}}_{u,m}^{n}\right]+\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}}\right)^{-1}\mathbf{y}_{u}^{\mathrm{C},j}\left(\mathbf{y}_{u}^{\mathrm{C},j}\right)^{\text{H}}\right]\right\},$
According to (47), the entropy matrix $h\left(\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)$ is obtained as in (48).
$h{\left(\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)}$ (48)
$=\sum_{n=0}^{N-1}\bigg{\{}MN_{\text{U}}{\log_{2}\pi+M}N_{\text{U}}{\log_{2}e}+M\log_{2}\left[\det\left(\sum_{b=1}^{\mathcal{B}}\left[\left(\tilde{\mathbf{H}}_{u,m}^{n}\right)^{\text{H}}E\left\{\left(\mathbf{X}_{n}^{b,j}\right)^{\text{H}}\mathbf{X}_{n}^{b,j}\right\}\mathbf{\tilde{H}}_{u,m}^{n}\right]+\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}}\right)\right]\bigg{\}}.$
Meanwhile, the noise entropy $h\left(\mathbf{W}_{u}^{\mathrm{C}}\right)$ can be obtained as
$h{\left(\mathbf{W}_{u}^{\mathrm{C}}\right)}$ $=NMN_{\mathrm{U}}\log_{2}\pi+NMN_{\mathrm{U}}\log_{2}e$ (49)
$\quad+NM\mathrm{log}_{2}{\left[\det{\left(\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}}\right)}\right]}.$
Therefore, substituting (48) and (49) into (45), the communication MI of the proposed CA-enabled MIMO-OFDM ISAC system in (46) is derived in (50),
$I\left(\mathbf{X}_{\mathrm{C}}^{1},\mathbf{X}_{\mathrm{C}}^{2};\mathbf{Y}_{u}^{\mathrm{C}}\mid\mathbf{H}_{\mathrm{C}}^{1},\mathbf{H}_{\mathrm{C}}^{2}\right)$ (50)
$=M\sum_{n=0}^{N-1}\log_{2}\left[\left(\sigma_{\mathrm{C}}^{2}\right)^{-N_{\text{U}}}\det\left(\sum_{b=1}^{\mathcal{B}}\left[\left(\tilde{\mathbf{H}}_{u,m}^{n}\right)^{\text{H}}E\left\{\left(\mathbf{X}_{n}^{b,j}\right)^{\mathrm{H}}\mathbf{X}_{n}^{b,j}\right\}\mathbf{\tilde{H}}_{u,m}^{n}\right]+\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}}\right)\right]$
$=M\log_{2}\left[\left(\sigma_{\mathrm{C}}^{2}\right)^{-N_{\text{U}}N}\prod_{n=0}^{N-1}\det\left(\sum_{b=1}^{\mathcal{B}}\left[\left(\tilde{\mathbf{H}}_{u,m}^{n}\right)^{\text{H}}E\left\{\left(\mathbf{X}_{n}^{b,j}\right)^{\text{H}}\mathbf{X}_{n}^{b,j}\right\}\mathbf{\tilde{H}}_{u,m}^{n}\right]+\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}}\right)\right]$
$=M\log_{2}\left[\left(\sigma_{\mathrm{C}}^{2}\right)^{-N_{\text{U}}N}\det\left(\sum_{b=1}^{\mathcal{B}}\left[\left(\mathbf{H}^{b}\right)^{\text{H}}\text{diag}\left\{E\left\{\left(\mathbf{X}_{\mathrm{C}}^{b,j}\right)^{\text{H}}\mathbf{X}_{\mathrm{C}}^{b,j}\right\}\right\}\mathbf{H}^{b}\right]+\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}N}\right)\right]$
$=M\log_{2}\left[\left(\sigma_{\mathrm{C}}^{2}\right)^{-N_{\text{U}}N}\det\left[\sum_{b=1}^{\mathcal{B}}\left[\left(\mathbf{H}^{b}\right)^{\text{H}}E\left\{\frac{\left(\mathbf{X}_{\mathrm{C}}^{b}\right)^{\text{H}}\mathbf{X}_{\mathrm{C}}^{b}}{M}\right\}\mathbf{H}^{b}\right]+\sigma_{\mathrm{C}}^{2}\mathbf{I}_{N_{\text{U}}N}\right)\right],$
where the $\mathbf{H}^{b}$ is the transformation of $\mathbf{H}_{\mathrm{C}}^{b}$ in (8), expressed as follows
$\mathbf{H}^{b}=\text{diag}\left\{\mathbf{\tilde{H}}_{u,m}^{0},\cdots,\mathbf{\tilde{H}}_{u,m}^{N-1}\right\}\in\mathbb{C}^{N_{\text{T}}N\times N_{\text{U}}N}.$ (51)
IV-B Sensing Performance Analysis
The CRLB is generally used to determine the lower bound on the variance of unbiased estimator and characterize the sensing ability of a signal [28]. Therefore, the CRLB of the CA-enabled MIMO-OFDM ISAC system is derived in this section.
IV-B1 CRLB of MIMO-OFDM system
The CRLB for OFDM ISAC system is derived in [7], and the CRLB for MIMO-OFDM ISAC system is revealed in Theorem 3. For ease of derivation, $\gamma=\frac{2v_{0}}{c}$ is used as an estimation of velocity to avoid the impact of different carrier frequency on theoretical derivation [7]. Meanwhile, $\sin{\theta_{\text{Rx}}}$ is used as an estimation of AoA to avoid the presence of unknown parameters after multiple derivations.
Theorem 3
For a MIMO-OFDM ISAC system (e.g., the $b$-th CC), the CRLBs of $\sin{\theta_{\text{Rx}}}$, $\tau_{0}$, and $\gamma$ are expressed as (52), (53), and (54), respectively.
$CRLB(\sin{\theta_{\text{Rx}}})=\frac{\Gamma_{\text{mimo}}N_{\text{U}}\left[M_{b}N_{b}\Theta_{b}^{n}\Theta_{b}^{m}-\left(\Theta_{b}^{m,n}\right)^{2}\right]}{M_{b}N_{b}},$ (52)
$CRLB(\tau_{0})=\frac{\Gamma_{\text{mimo}}N_{b}\left[M_{b}N_{\text{U}}\Theta_{b}^{p}\Theta_{b}^{m}-\left(\Theta_{b}^{p,m}\right)^{2}\right]}{M_{b}N_{\text{U}}},$ (53)
$CRLB(\gamma)=\frac{\Gamma_{\text{mimo}}M_{b}\left[N_{b}N_{\text{U}}\Theta_{b}^{p}\Theta_{b}^{n}-\left(\Theta_{b}^{p,n}\right)^{2}\right]}{N_{b}N_{\text{U}}},$ (54)
where $\Gamma_{\text{mimo}}$ is denoted by (55),
$\Gamma_{\text{mimo}}=\frac{1}{{\varpi\cdot\left(2\pi\right)^{2}}}\frac{1}{M_{b}N_{b}N_{\text{U}}\Theta_{b}^{p}\Theta_{b}^{n}\Theta_{b}^{m}+2\Theta_{b}^{p,n}\Theta_{b}^{p,m}\Theta_{b}^{m,n}-N_{b}\left(\Theta_{b}^{p,m}\right)^{2}\Theta_{b}^{n}-M_{b}\left(\Theta_{b}^{p,n}\right)^{2}\Theta_{b}^{m}-N_{\text{U}}\left(\Theta_{b}^{m,n}\right)^{2}\Theta_{b}^{p}},$ (55)
$\varpi=\frac{\left|A\right|^{2}}{\sigma^{2}}$, $\Theta_{b}^{n}=\sum\limits_{n}\left(n\Delta f^{b}\right)^{2}$, $\Theta_{b}^{m}=\sum\limits_{m}\left(mf_{\mathrm{C}}^{b}T^{b}\right)^{2}$, $\Theta_{b}^{p}=\sum\limits_{p}\left(p\frac{d_{\text{r}}}{\lambda^{b}}\right)^{2}$, $\Theta_{b}^{m,n}=\sum\limits_{m}\sum\limits_{n}mn\Delta f^{b}f_{\mathrm{C}}^{b}T^{b}$, $\Theta_{b}^{p,n}=\sum\limits_{p}\sum\limits_{n}pn\frac{d_{\text{r}}}{\lambda^{b}}\Delta f^{b}$, and $\Theta_{b}^{p,m}=\sum\limits_{p}\sum\limits_{m}pm\frac{d_{\text{r}}}{\lambda^{b}}f_{\mathrm{C}}^{b}T^{b}$.
Proof:
Please refer to Appendix A. ∎
IV-B2 CRLB of CA-enabled MIMO-OFDM system
The CRLBs of range and velocity estimations of the proposed ISAC system are derived. For simplicity, two CCs occupy the same number of resource elements (REs), i.e., $N_{1}=N_{2}=N$ and $M_{1}=M_{2}=M$.
According to (12), the CRLBs of $\tau_{0}$ and $\gamma$ for the proposed CA-enabled MIMO-OFDM ISAC system are expressed as (56) and (57), respectively.
$CRLB_{\text{CA}}(\tau_{0})=\frac{\Gamma_{\text{ca}}N\left[MN_{\text{U}}\Theta_{B}^{p}\Theta_{B}^{m}-\left(\Theta_{B}^{p,m}\right)^{2}\right]}{MN_{\text{U}}},$ (56)
$CRLB_{\text{CA}}(\gamma)=\frac{\Gamma_{\text{ca}}M\left[NN_{\text{U}}\Theta_{B}^{p}\Theta_{B}^{n}-\left(\Theta_{B}^{p,n}\right)^{2}\right]}{NN_{\text{U}}},$ (57)
$\Gamma_{\text{ca}}=\frac{1}{\varpi\cdot\left(2\pi\right)^{2}}\frac{1}{MNN_{\text{U}}\Theta_{B}^{p}\Theta_{B}^{n}\Theta_{B}^{m}+2\Theta_{B}^{p,n}\Theta_{B}^{p,m}\Theta_{B}^{m,n}-N\left(\Theta_{B}^{p,m}\right)^{2}\Theta_{B}^{n}-M\left(\Theta_{B}^{p,n}\right)^{2}\Theta_{B}^{m}-N_{\text{U}}\left(\Theta_{B}^{m,n}\right)^{2}\Theta_{B}^{p}},$ (58)
where $\Gamma_{\text{ca}}$ is expressed in (58), $\Theta_{B}^{n}=\sum\limits_{n}\left(n\left(\Delta f^{1}+\Delta f^{2}\right)\right)^{2}$, $\Theta_{B}^{m}=\sum\limits_{m}\left(m\left(f_{\mathrm{C}}^{1}T^{1}+f_{\mathrm{C}}^{2}T^{2}\right)\right)^{2}$, $\Theta_{B}^{p}=\sum\limits_{p}\left(p\left(\frac{d_{\text{r}}}{\lambda^{1}}+\frac{d_{\text{r}}}{\lambda^{2}}\right)\right)^{2}$, and
$\Theta_{B}^{m,n}=\sum\limits_{m}\sum\limits_{n}mn\left(\Delta f^{1}+\Delta f^{2}\right)\left(f_{\mathrm{C}}^{1}T^{1}+f_{\mathrm{C}}^{2}T^{2}\right),$
$\Theta_{B}^{p,n}=\sum\limits_{p}\sum\limits_{n}pn\left(\frac{d_{\text{r}}}{\lambda^{1}}+\frac{d_{\text{r}}}{\lambda^{2}}\right)\left(\Delta f^{1}+\Delta f^{2}\right),$
$\Theta_{B}^{p,m}=\sum\limits_{p}\sum\limits_{m}pm\left(\frac{d_{\text{r}}}{\lambda^{1}}+\frac{d_{\text{r}}}{\lambda^{2}}\right)\left(f_{\mathrm{C}}^{1}T^{1}+f_{\mathrm{C}}^{2}T^{2}\right).$
Substituting $\Delta f^{2}=Q\Delta f^{1}$ and $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$ mentioned in Section III-B into (56) and (57), the simplest close-form expressions for CRLBs of $\tau_{0}$ and $\gamma$ are
$CRLB_{\text{CA}}(\tau_{0})=\frac{\Gamma_{\text{CA}}\left[\frac{(2N_{\text{U}})(2M-1)}{36}-\frac{(N_{\text{U}}-1)(M-1)}{16}\right]}{\left(\Delta f^{1}\right)^{2}\left(Q+1\right)^{2}\left(N-1\right)},$ (59)
$CRLB_{\text{CA}}(\gamma)=\frac{\Gamma_{\text{CA}}\left[\frac{(2N_{\text{U}})(2N-1)}{36}-\frac{(N_{\text{U}}-1)(N-1)}{16}\right]}{\left(2f_{\mathrm{C}}^{1}T^{1}\right)^{2}\left(M-1\right)},$ (60)
where $\Gamma_{\text{CA}}$ is expressed in (61).
$\Gamma_{\text{CA}}=\frac{1}{\varpi(2\pi)^{2}MNN_{\text{U}}}\times$ (61)
$\frac{1}{\frac{(2N_{\text{U}}-1)}{6}\frac{(2N-1)}{6}\frac{(2M-1)}{6}+(N_{\text{U}}-1)\left[\frac{(N-1)(M-1)}{32}-\frac{(M-1)(2N-1)}{96}-\frac{(N-1)(2M-1)}{96}\right]-\left(\frac{(M-1)(N-1)}{16}\right)\left(\frac{(2N_{\text{U}}-1)}{6}\right)}.$
According to $\tau_{0}=\frac{2R}{c}$ and $\gamma=\frac{2v_{0}}{c}$, the CRLBs of $R$ and $v_{0}$ of CA-enabled MIMO-OFDM ISAC system are
$CRLB_{\text{CA}}(R)=\frac{c^{2}}{4}\cdot\frac{\Gamma_{\text{CA}}\left[\frac{(2N_{\text{U}})(2M-1)}{36}-\frac{(N_{\text{U}}-1)(M-1)}{16}\right]}{\left(\Delta f^{1}\right)^{2}\left(Q+1\right)^{2}\left(N-1\right)},$ (62)
$CRLB_{\text{CA}}(v_{0})=\frac{c^{2}}{4}\cdot\frac{\Gamma_{\text{CA}}\left[\frac{(2N_{\text{U}})(2N-1)}{36}-\frac{(N_{\text{U}}-1)(N-1)}{16}\right]}{\left(2f_{\mathrm{C}}^{1}T^{1}\right)^{2}\left(M-1\right)}.$ (63)
V Simulation Results and Analysis
In this section, the simulation results are provided to verify the high-spend communication and high-accuracy sensing capabilities of the proposed CA-enabled MIMO-OFDM ISAC system. Simulation parameters are detailed in Table II. The simulation results are obtained with 1000 times Monte Carlo simulations.
TABLE II: Simulation parameters [7, 29, 30].
Symbol Parameter Value Symbol Parameter Value
$f_{\mathrm{C}}^{1}$ Carrier frequency of low-frequency band 5.9 GHz [7] $f_{\mathrm{C}}^{2}$ Carrier frequency of high-frequency band 24 GHz [7]
$M_{1}$ Number of OFDM symbols on low-frequency band 64 $M_{2}$ Number of OFDM symbols on high-frequency band 64
$N_{1}$ Number of subcarriers on low-frequency band 512 $N_{2}$ Number of subcarriers on high-frequency band 512
$T^{1}$ Total symbol duration on low-frequency band 39.5 $\mathrm{\mu}$s $T^{2}$ Total symbol duration on high-frequency band 9.7 $\mathrm{\mu}$s
$\Delta f^{1}$ Subcarrier spacing on low-frequency band 30 kHz [30] $\Delta f^{2}$ Subcarrier spacing on high-frequency band 120 kHz [30]
$N_{\text{T}}$ Number of transmit antenna of BS 4 $N_{\text{R}}$ Number of received antenna of BS 4
$N_{\text{U}}$ Number of received antenna of UE {4, 5, 6} $\theta_{\text{Rx}},\theta_{\text{Tx}}$ AoA and AoD of target $30^{\circ}$
$r_{0}$ Range of target 117 m $v_{0}$ Velocity of target 13 m/s
V-A Performance of Communication
According to (43) and (46) derived in Section IV-A, the communication MIs of CA-enabled MIMO-OFDM ISAC system and MIMO-OFDM ISAC system under various SNR and $N_{\text{U}}$ are simulated in Fig.  5.
Figure 5: Communication MI of the proposed CA-enabled MIMO-OFDM ISAC system and MIMO-OFDM ISAC system.
In Fig. 5, the communication MI of the CA-enabled MIMO-OFDM ISAC signal is larger than that of the MIMO-OFDM ISAC signal under different $N_{\text{U}}$.
V-B Performance of Sensing
In this section, the CRLBs and RMSEs of range and velocity estimations are simulated to demonstrate the high-accuracy sensing performance of the proposed CA-enabled MIMO-OFDM ISAC system.
V-B1 CRLB
According to the closed-form sensing CRLBs derived in Section IV-B, the CRLBs for range and velocity estimations of CA-enabled MIMO-OFDM ISAC signal, MIMO-OFDM ISAC signal, and OFDM ISAC signal are simulated, as shown in Figs. 6 and 7. Fig. 6 reveals the following phenomena.
  • The CRLB of the CA-enabled MIMO-OFDM ISAC signal is the lowest, indicating that the proposed CA-enabled MIMO-OFDM ISAC signal has better range estimation performance than MIMO-OFDM ISAC signal and OFDM ISAC signal.
  • Comparing with the CRLB of low-frequency MIMO-OFDM ISAC signal, the CRLB of high-frequency MIMO-OFDM ISAC signal is lower, which is caused by the large bandwidth resources of high-frequency signal.
  • Comparing with the CRLB of OFDM ISAC signal, the CRLB of MIMO-OFDM ISAC signal is lower, which benefits from the spatial diversity of MIMO.
Fig. 7 demonstrates the following phenomena.
  • The CRLB of CA-enabled MIMO-OFDM ISAC signal is lowest, indicating that the CA-enabled MIMO-OFDM ISAC signal has better performance of velocity estimation than MIMO-OFDM ISAC signal and OFDM ISAC signal.
  • Upon observing Fig. 7(b), the high-frequency MIMO-OFDM ISAC signal and low-frequency MIMO-OFDM ISAC signal are overlapped, which also exist in the OFDM ISAC signals. The reason is as follows. According to (63), for the ISAC signals occupying the same number of REs and satisfying $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$ in simulation parameters, the CRLBs of velocity estimation are identical.
  • Comparing with the CRLB of OFDM ISAC signal, the CRLB of MIMO-OFDM ISAC signal is lower, which also benefits from the spatial diversity of MIMO.
Therefore, the simulation results indicate the superior sensing performance of the proposed CA-enabled MIMO-OFDM ISAC signal compared with MIMO-OFDM ISAC signal and OFDM ISAC signal.
Figure 6: Comparison of CRLB for range estimation
(a) 2D view
(b) 3D view
Figure 7: Comparison of CRLB for velocity estimation
V-B2 RMSE
The data-level and signal-level fusion methods are compared in Fig. 8, where the the following phenomena are revealed.
  • In Fig. 8, the RMSE of symbol-level fusion method is consistently lower than that of the data-level fusion method, indicating the superiority of the proposed symbol-level fusion method over the data-level fusion method.
  • Compared with the data-level fusion method, the accuracy of range and velocity estimations under symbol-level fusion method is improved by up to 66.91% and 95.67%, respectively.
Then, the RMSEs for range and velocity estimations of the proposed CA-enabled MIMO-OFDM ISAC signal and MIMO-OFDM ISAC signal are simulated in Fig. 9. There are several assumptions.
  1. 1.
    The SNR in the simulation is the transmit SNR of single RE.
  2. 2.
    The theoretical RMSEs of range and velocity estimations are expressed in (64) and (65), respectively [31].
    $\sigma_{R}=\frac{c}{2B\sqrt{2\chi}},$ (64)
    $\sigma_{V}=\frac{c}{2M_{\text{ofdm}}f_{\mathrm{C}}T_{\mathrm{C}}\sqrt{2\chi}},$ (65)
    where $B$ and $\chi$ are the bandwidth and received SNR of signal, respectively; $M_{\text{ofdm}}$ is the number of OFDM symbols and $f_{\mathrm{C}}$ is the carrier frequency; $T_{\mathrm{C}}$ is the total symbol duration.
(a) RMSE for range estimation
(b) RMSE for velocity estimation
Figure 8: Comparison of data-level fusion and symbol-level sensing data fusion methods
Fig. 9(a) reveals the following phenomena.
  • With the increase of SNR, the RMSEs of CA-enabled MIMO-OFDM ISAC signal and MIMO-OFDM ISAC signal are stable, because the accuracy of sensing method is limited by the grid size.
  • Comparing with the RMSEs of MIMO-OFDM ISAC signals, the RMSE of CA-enabled MIMO-OFDM ISAC signal is lower, indicating the superior performance of range estimation of the proposed CA-enabled MIMO-OFDM ISAC signal over the MIMO-OFDM ISAC signal.
  • When SNR is smaller than $-20$ dB, comparing with the RMSE of low-frequency MIMO-OFDM ISAC signal, the RMSE of high-frequency MIMO-OFDM ISAC signal is higher. In the simulation, we employed a typical path loss model and set the received SNR of high-frequency signal 5 dB lower than that of low-frequency signal [32]. Therefore, for comparison under the same received SNR, the curve of high-frequency MIMO-OFDM ISAC signal needs to shift 5 dB to the left.
According to (65), since $f_{\mathrm{C}}^{1}T^{1}=f_{\mathrm{C}}^{2}T^{2}$ in the simulation, the differences in the RMSEs of high and low frequency signals are only the received SNR.
(a) RMSE for range estimation
(b) RMSE for velocity estimation
Figure 9: Comparison of MIMO-OFDM ISAC signal on different frequency bands and CA-enabled MIMO-OFDM ISAC signal.
VI Conclusion
In this paper, CA-enabled MIMO-OFDM ISAC system is investigated, where CA aggregates fragmented high and low-frequency bands, achieving the high-speed communication and high-accuracy sensing. Firstly, we introduce the signal model of the proposed CA-enabled MIMO-OFDM ISAC system and the ISAC signal processing framework. To address the challenges in the sensing signal processing introduced by CA, symbol-level fusion method of the sensing data on high and low-frequency bands is proposed. In the proposed symbol-level fusion method, the phase alignment of the echo signals on high and low-frequency bands is realized by the angle compensation, space-domain diversity and vector cross-correlation operations. Then, the fusion algorithms of feature vectors are proposed to realize high-accuracy sensing. The simulation results reveal that the performance of symbol-level fusion method is better than that of data-level fusion method. Furthermore, the closed-form communication MI and sensing CRLB of the proposed ISAC signal are derived. Numerical results reveal that the proposed CA-enabled MIMO-OFDM ISAC signal exhibits superior communication rate and high-accuracy sensing performance over the conventional MIMO-OFDM ISAC signal. This work reveals the potential of CA-enabled MIMO-OFDM ISAC signal in enhancing communication and sensing capabilities.
Appendix A proof of theorem 3
According to (12), the received echo signal by the $p$-th received antenna on the $n$-th subcarrier during the $m$-th OFDM symbol time is expressed as
$y_{p,m,n}=$ $Ae^{j2\pi mf_{\mathrm{C}}^{b}T^{b}\gamma}e^{-j2\pi n\Delta f^{b}\tau_{0}}e^{j2\pi p\left(\frac{d_{\text{r}}}{\lambda^{b}}\right){\sin\theta_{\text{Rx}}}}$ (66)
$+w_{p,m,n},$
where $A\in\mathbb{C}$ and $w_{p,m,n}\sim\mathcal{CN}(0,\sigma^{2})$ is the AWGN.
The likelihood function for the joint estimation of $\sin\theta_{\text{Rx}}$, $\tau_{0}$, and $\gamma$ is
$\ln f\left(y;\sin\theta_{\text{Rx}},\tau_{0},\gamma\right)$ (67)
$=-\frac{M_{b}N_{b}N_{\text{U}}}{2}\ln\left(2\pi\sigma^{2}\right)$
$\quad-\frac{1}{2\sigma^{2}}\sum_{p}\sum_{m}\sum_{n}\left(y_{p,m,n}-s_{p,m,n}\right)^{*}\left(y_{p,m,n}-s_{p,m,n}\right),$
where the PDF is
$f\left(y;\sin\theta_{\text{Rx}},\tau_{0},\gamma\right)$ (68)
$=\frac{1}{\left(2\pi\sigma^{2}\right)^{M_{b}N_{b}N_{\text{U}}/2}}e^{-\frac{1}{2\sigma^{2}}\sum\limits_{p}\sum\limits_{m}\sum\limits_{n}\left|y_{p,m,n}-s_{p,m,n}\right|^{2}},$
with
$s_{p,m,n}=Ae^{j2\pi mf_{\mathrm{C}}^{b}T^{b}\gamma}e^{-j2\pi n\Delta f^{b}\tau_{0}}e^{j2\pi p\left(\frac{d_{\text{r}}}{\lambda^{b}}\right){\sin\theta_{\text{Rx}}}}.$ (69)
Based on the relationship between the Fisher information matrix (FIM) and CRLB, we can obtain that
$\mathbf{F}^{-1}$ $=\left[\begin{array}[]{ccc}F_{\sin{\theta_{\text{Rx}}},\sin{\theta_{\text{Rx}}}}&F_{\sin{\theta_{\text{Rx}}},\tau_{0}}&F_{\sin{\theta_{\text{Rx}}},\gamma}\\ F_{\tau_{0},\sin{\theta_{\text{Rx}}}}&F_{\tau_{0},\tau_{0}}&F_{\tau_{0},\gamma}\\ F_{\gamma,\sin{\theta_{\text{Rx}}}}&F_{\gamma,\tau_{0}}&F_{\gamma,\gamma}\end{array}\right]^{-1}$ (70)
$=\left[\begin{array}[]{ccc}CRLB(\sin{\theta_{\text{Rx}}})&Cov(\sin{\theta_{\text{Rx}}},\tau_{0})&Cov(\sin{\theta_{\text{Rx}}},\gamma)\\ Cov(\tau_{0},\sin{\theta_{\text{Rx}}})&CRLB(\tau_{0})&Cov(\tau_{0},\gamma)\\ Cov(\gamma,\sin{\theta_{\text{Rx}}})&Cov(\gamma,\tau_{0})&CRLB(\gamma)\end{array}\right],$
where $Cov(\cdot)$ is the covariance operation and $F_{\alpha,\beta}=-E\left[\frac{\partial^{2}\ln f\left(x;\alpha,\beta,...\right)}{\partial\alpha\partial\beta}\right]$.
According to (67) and (70), the CRLBs of angle, range, and velocity estimations for MIMO-OFDM ISAC system are derived, as shown in Theorem 3.